OpenStudy (anonymous):

11. If the mass of a proton is 2.53 x 10-37 gram, what is the mass of 1,000 protons

OpenStudy (zehanz):

It will be 1000 times more, or 10³ times more, so:\[2.53 \cdot 10^{-37} \cdot 10^3\]Now use the exponent rule:\[a^b \cdot a^c=a^{b+c}\]to get that nice-looking, standard notation answer!